OpenStudy (anonymous):

Evaluate. 16 • 3 + [(17 – 3) ÷ 2] – 43

OpenStudy (gorv):

first solve what is inside the bracket

OpenStudy (gorv):

bracket is solved first

OpenStudy (gorv):

17-3=????

OpenStudy (anonymous):

14

OpenStudy (gorv):

@lovergirl09876

OpenStudy (gorv):

\[14 \div 2 =???\]

OpenStudy (gorv):

@lovergirl09876

OpenStudy (anonymous):

7

OpenStudy (gorv):

now we are done with bracket

OpenStudy (gorv):

\[16*3+7-43\] we left with this now

OpenStudy (gorv):

multiplication has high order so do it first

OpenStudy (gorv):

16*3=???

OpenStudy (anonymous):

48

OpenStudy (gorv):

48+7-43

OpenStudy (gorv):

now addition is done first 48+7=???

OpenStudy (anonymous):

55

OpenStudy (gorv):

55-43=??

OpenStudy (anonymous):

12

OpenStudy (gorv):

you got it right .....problem solved :)

OpenStudy (anonymous):

thanks for your help

OpenStudy (gorv):

welcome :)